The demand function of a good is given by the expression Q³-8Q²+20Q+2
I) Find the level of output, which maximizes total revenue
II) Find the maximum profit
1
Expert's answer
2020-05-24T18:07:36-0400
"D(Q)=Q^3-8Q^2+20Q+2"

i)


"3Q^2-16Q+20=0"


"Q_{max}=2"

ii)


"p=2^2-8\times2^2+20\times2+2=18"


"Pr=18\times2=36"